Circular No. 7271

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999ed IN UGC 3555 M. Modjaz and W. D. Li, University of California at Berkeley (UCB),

on behalf of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126), report their discovery of an apparent supernova (mag ∼ 17.8) on Oct. 5.5 UT with the 0.8-m Katzman Automatic Imaging Telescope (KAIT). SN 1999ed is located at α = 6^{h}50^{m}00^{s}.95, δ = +25^{o}37′54′′.5 (equinox 2000.0), which is 17′′.4 east and 9′′.0 south of the nucleus of UGC 3555. A KAIT image taken on 1998 Nov. 21 showed nothing at the position of SN 1999ed (limiting mag 19.5).

Q. Y. Qiao, Y. L. Qiu, and J. Y. Hu, Beijing Astronomical Observatory (BAO), report the independent discovery of this object during the course of the BAO Supernova Survey. It was found (mag ∼ 18.5) in an unfiltered CCD image taken on Oct. 6.88 UT with the BAO 0.6-m telescope, and confirmed in an image taken on Oct. 7.83. CCD frames taken on Mar. 28.53 (limiting mag ∼ 19.0) showed nothing at the position of the supernova.

A. V. Filippenko and R. T. Chornock, UCB, report that CCD spectra (range 330–1000 nm) obtained on Oct. 8 UT with the Shane 3-m reflector at Lick Observatory reveal that the object is a normal type-II supernova, probably < 2 months after the explosion; there are strong Balmer lines with P-Cyg profiles. The redshift, measured from narrow emission lines in the host galaxy, is 0.016.

GM SAGITTARII M. R. Garcia and J. E. McClintock, Harvard-Smithsonian Center for

Astrophysics; and P. J. Callanan, University College, Cork, report: “We have obtained a total of 25 spectra (range 380-750 nm, resolution 0.6 nm) with the Whipple Observatory 1.52-m reflector (+ FAST spectrograph) of GM Sgr, centered on the nights of Sept. 17.2, 18.2, and 19.1 UT. During this time, the Hα line transitioned from emission to absorption. Equivalent widths were −7(1), −3(1), and +0.45(5) nm, respectively. The strong emission line seen on Sept. 17.2 had FWHM = 3.3(3) nm and a complex structure with two prominent peaks on the blue wing. The remainder of the Balmer series lines (through H_{16}) are visible in each of the spectra; they are more narrow (1.0 nm < FWHM < 2.0 nm) than the Hα emission line and were always in absorption. The continuum and absorption-line spectrum resembles that of an A star. The velocities of the Balmer absorption lines measured on the three consecutive nights were +320(20), 0(40), and 80(20) km/s, respectively.”

1999 October 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green